# Learning Molecular Representation in a Cell

### v2

@@ 4.3 Theoretical Motivation for Decoder-based Alignment
InfoNCE [35] is the contrastive loss used for most CLIP-like methods [39, 53]. In this work, we show that the mutual information lower bound based on InfoAlign is tighter than that based on InfoNCE.
@@
For the molecular representation $Z$ and any of cell morphology, gene expressions, and molecular fingerprints $Y$, the encoder-based mutual information lower bound $I_{ELB}$ for InfoNCE can be derived by incorporating $K-1$ additional samples from the modality of $Y$ (denoted as $Y_{2:K}$) to build the Monte Carlo estimate $m(\cdot)$ of the partition function, as follows [34, 37]:
@@ 5 Experiments
5.1 Implementation of Context Graph and Pretraining Setting
Data Source of Context Graph
We create the context graph based on (1) two Cell Painting datasets [4, 6], containing around 140K molecule perturbations (molecule and cell morphology pairs) and 15K genetic perturbations (gene and cell morphology pairs) across 1.6 billion human cells; (2) Hetionet [16], which captures gene-gene and gene-molecule relationships from millions of biomedical studies; and (3) a dataset reporting differential gene expression values for 978 landmark genes [54] for chemical perturbations (molecule and gene expression pairs) [49].
Node Features
Different profiling methods provide node features in different ways. Morgan fingerprints [42] are feature vectors extracted from each molecule’s structure, CellProfiler features [5] are computed from the image of each cell and represent cell morphology, and L1000 profiles [49] capture gene expression values on 978 landmark genes from cells treated with a chemical perturbation. Here are two practical considerations for the context graphs: (1)  Chandrasekaran et al. [6] provided one dataset that measured the cell morphology impacts of perturbing individual genes. The 15K genetic perturbations [6] provide gene-cell morphology pairs but lack corresponding gene expression profiles. Still, we keep the gene nodes from this dataset to account for potential gene-gene interactions and incorporate cell morphology features into them. (2) All 978 landmark genes have expression values linked to the molecules used in [54]. We update new gene expression nodes with 978-dimensional feature vectors. These vectors summarize all molecule-gene expression connections for a small molecule perturbation. This approach efficiently reduces dense connections between landmark genes and molecules. We select the top 1% of gene-molecule expression values as new edges to enrich the context graph’s connectivity. We scale cell morphology and gene expression features to a range of 0 to 1 using the Min-Max scaler along each dimension.
Edge Weights
We assign a weight of 1 to edges defined by chemical perturbations from [4, 6]. For edges defined by computational criteria, we compute cosine similarity for nodes if they are in the same feature space (such as two cell morphology/gene expression profiles, or Morgan fingerprints). To avoid noisy edges from computations, we (1) apply a 0.8 threshold for cosine similarity, and additionally (2) explicitly enforce 99.5% sparsity by selecting top similar edges.
Encoder and Decoder
We use the Graph Isomorphism Model (GIN) [56] as the molecular structure encoder. All molecules on the context graph are used to pretrain the molecular representation encoder. Because we extract feature vectors as the decoding targets in different modalities, we could efficiently use Multi-Layer Perception (MLP) as modality decoders. In each training batch, random walks start from the molecule node to extract the walk path. Then, decoders are pretrained to reconstruct corresponding node features from nodes over the path. More details are in appendix B.
Next, we focus on three research questions (RQs) regarding InfoAlign’s representation for molecular property prediction, molecule-morphology matching, and hyperparameter analysis.
5.2 RQ1: Molecular Property Prediction
@@ 5.2.1 Experimental Setting
We fine-tune MLPs on various representation learning approaches for predicting molecular properties.
@@
Dataset and Evaluation. We select datasets for important tasks in drug discovery, including activity classification for various assays in ChEMBL2K [13] and Broad6K [32], drug toxicity classification using ToxCast [41], and absorption, distribution, metabolism, and excretion (ADME) regression using Biogen3K [11]. The dataset statistics are in Table 1. We apply scaffold-splitting for all datasets. We follow [18] for the ToxCast dataset, and a 0.6:0.15:0.25 ratio for training, validation, and test sets for other datasets. We use the Area under the curve (AUC) for classification and mean absolute error (MAE) for regression. Mean and standard deviations are reported from ten runs.
@@
Baseline. We consider baselines from three representation sources: molecular structures, cell morphology, and gene expression. These expand into six categories: (1) Molecular descriptors/fingerprints (FP) of chemical structure [42]; (2) pretrained GNN representations of chemical structure from {AttrMask, ContextPred, EdgePred} [17]; (3) pretrained chemical language model (ChemLM) representations of chemical structure from {Roberta, GPT2, MolT5, ChemGPT} [12]; (4,5) cell morphology and differential gene expression values from cells treated with each molecule [54, 6]; (6) CLOOME [45] and InfoCORE [53] for multi-modal alignment, using all three data types (structure, morphology, and gene expression). We use MLPs, Random Forests (RF), and Gaussian Processes (GP) for representations for (1,4,5). We implemented up to 19 baselines, presenting the best-performing result in each of categories (1)-(5). Setting details and all results are in Sections C.1 and C.3.
Table 2: Results on ChEMBL2K and Broad6K. We report average AUC (Avg.), as well as the percentage of tasks achieving AUC above 80%, 85%, and 90%. Best and second best mean are bold and underline, respectively.
ChEMBL2K (AUC $\uparrow$) Broad6K (AUC $\uparrow$) (# Molecule / # Task) (2355 / 41) (6567 / 32) Method Avg. > 80 % > 85 % > 90 % Avg. > 80 % > 85 % > 90 % Morgan FP 76.8±2.2 48.8±3.9 34.6±6.3 21.9±5.7 63.3±0.3 6.3±0.0 4.4±1.7 3.1±0.0 Pretrained GNN 77.0±0.5 55.1±1.3 34.2±4.6 14.6±1.7 59.9±0.2 7.5±1.7 3.1±0.0 3.1±0.0 Pretrained ChemLM 74.7±1.9 46.3±3.4 35.1±4.4 22.9±1.3 60.6±0.3 7.5±1.7 3.1±0.0 1.9±1.7 Cell Morphology 64.3±2.4 15.6±6.6 8.3±3.7 4.9±3.9 55.3±0.1 0.0±0.0 0.0±0.0 0.0±0.0 Gene Expression 56.1±1.1 5.1±1.4 3.4±1.3 3.4±1.3 56.9±1.4 1.9±1.7 1.9±1.7 1.9±1.7 CLOOME 66.7±1.8 26.8±4.6 16.1±3.7 10.7±5.1 61.7±0.4 3.1±0.0 3.1±0.0 0.0±0.0 InfoCORE (GE) 79.3±0.9 62.4±2.8 46.3±3.0 30.3±2.2 60.2±0.2 3.1±0.0 0.0±0.0 0.0±0.0 InfoCORE (CP) 73.8±2.0 37.6±9.2 26.3±4.7 10.7±4.1 61.1±0.2 6.3±0.0 3.1±0.0 0.0±0.0 InfoAlign (Ours) 81.3±0.6 66.3±2.7 49.3±2.7 35.1±3.7 70.0±0.1 18.8±2.2 3.1±0.0 3.1±0.0
@@
Table 3: Results on ToxCast and Biogen3K. We report the average AUC and the percentage of AUC above 80% on ToxCast, and regression MAE (scaled by × 100) for Biogen3K. Best and second best mean are bold and underline, respectively.
ToxCast (AUC $\uparrow$) Biogen3K (MAE ×100$\times$ 100$\downarrow$ 100 ↓ ) (# Molecule / # Task) (8576 / 617) (3521 / 6) Method Avg. >80 % Avg. hPPB rPPB RLM HLM ER Solubility Morgan FP 57.6±1.0 1.6±0.3 52.8±0.2 44.2±0.1 44.2±0.1 42.0±0.2 67.7±0.7 66.9±0.9 51.6±0.1 Pretrained GNN 63.5±1.1 4.8±3.0 67.3±0.3 82.4±1.1 49.8±0.7 51.7±1.0 57.9±0.6 62.6±0.5 99.1±1.2 Pretrained ChemLM 64.7±0.9 3.6±1.1 65.1±0.5 76.7±2.1 55.9±1.1 49.2±1.0 70.3±0.8 73.1±1.0 65.3±1.7 CLOOME 54.2±0.9 0.9±0.2 64.3±0.4 65.2±1.5 56.9±0.8 44.2±0.8 70.7±0.4 73.6±0.8 75.0±2.1 InfoCORE (GE) 65.3±0.2 5.4±1.7 69.9±1.2 79.9±3.6 51.6±1.8 51.3±2.1 78.6±0.3 77.8±1.9 80.3±0.9 InfoCORE (CP) 62.4±0.4 1.3±0.5 71.0±0.6 74.5±4.9 53.5±0.7 53.6±2.1 80.8±1.5 79.4±3.4 84.4±1.0 InfoAlign (Ours) 66.4±1.1 6.6±1.6 49.4±0.2 39.7±0.4 39.2±0.3 40.5±0.6 66.7±1.7 62.0±1.5 48.4±0.6
5.2.2 Results and Analysis
Results for molecular property prediction across various assays are presented in Tables 2 and 3 and Figure 3. Results for all 19 baselines are in Section C.3. Key observations include:
(1) Molecular structures are superior compared to cell morphology and gene expression features for predicting various molecular assays. This is likely because the datasets and tasks we selected fundamentally involve predicting the binding affinity of a molecule to a protein [13]; furthermore, in these datasets, molecules with activity in a given assay tend to have highly related structures, rather than representing two or more structurally distinct classes of molecules with activity; together this implies that molecular structure alone will tend to yield strong results. When comparing the three popular structure-based representation approaches, fingerprints outperform the others on Broad6K and Biogen3K. Pretrained GNN and ChemLM representations excel on ChEMBL2K and ToxCast, respectively. These findings suggest that continued efforts in better structure-based representation learning are still necessary.
(2) Cell morphology and gene expression features may complement molecular structures, yielding more generalizable representations. As shown in Figure 3, cell morphology and gene expression outperform molecular structure in approximately 20% and 36% of tasks on the ChEMBL2K and Broad6K datasets, respectively. This suggests that incorporating cell context into representation learning would be beneficial. That said, existing multi-modal baselines (InfoCORE, CLOOME) only outperform molecular structure-based approaches on ChEMBL2K and ToxCast, as they do not construct molecular representations holistically by using all cell-related modalities.
(3) InfoAlign achieves the best performance by capturing bottleneck representation across molecules, cell morphology, and gene expression. InfoAlign improves classification and regression performance by +10.58% and +6.33% compared to the best baselines on the Broad6K and Biogen3K datasets, respectively. This improvement is even more dramatic when using the 80% AUC threshold on classification datasets. While InfoCORE (GE) is the best among baselines for the ChEMBL2K and ToxCast datasets, it cannot align the molecular representation with more than two modalities and sometimes leads to negative transfer as observed in Broad6K and Biogen3K.

### v3

@@ 4.3 Theoretical Motivation for Decoder-based Alignment
InfoNCE [36] is the contrastive loss used for most CLIP-like methods [40, 54]. In this work, we show that the MI lower bound based on InfoAlign is tighter than that based on InfoNCE.
@@
For the molecular representation $Z$ and target $Y$ (from cell morphology, gene expressions, or molecular fingerprints), the encoder-based MI lower bound $I_{ELB}$ for InfoNCE can be derived by incorporating $K-1$ additional samples, denoted as $y_{2:K}$, to build the Monte Carlo estimate $m(\cdot)$ of the partition function [35, 38]:
@@
5 Implementation of Context Graph and Pretraining Setting
Data Source of Context Graph: We create the context graph based on (1) two Cell Painting datasets [4, 6], containing around 140K molecule perturbations (molecule and cell morphology pairs) and 15K genetic perturbations (gene and cell morphology pairs) across 1.6 billion human cells; (2) Hetionet [16], which captures gene-gene and gene-molecule relationships from millions of biomedical studies; and (3) a dataset reporting differential gene expression values for 978 landmark genes [56] for chemical perturbations (molecule and gene expression pairs) [50].
Node Features: Different profiling methods provide node features in different ways. Morgan fingerprints [43] are feature vectors extracted from each molecule’s structure, CellProfiler features [5] are computed from the image of each cell and represent cell morphology, and L1000 profiles [50] capture gene expression values on 978 landmark genes from cells treated with a chemical perturbation. Here are two practical considerations for the context graphs: (1)  Chandrasekaran et al. [6] provided one dataset that measured the cell morphology impacts of perturbing individual genes. The 15K genetic perturbations [6] provide gene-cell morphology pairs but lack corresponding gene expression profiles. Still, we keep the gene nodes from this dataset to account for potential gene-gene interactions and incorporate cell morphology features into them. (2) All 978 landmark genes have expression values linked to the molecules used in [56]. We update new gene expression nodes with 978-dimensional feature vectors. These vectors summarize all molecule-gene expression connections for a small molecule perturbation. This approach efficiently reduces dense connections between landmark genes and molecules. We select the top 1% of gene-molecule expression values as new edges to enrich the context graph’s connectivity. We scale cell morphology and gene expression features to a range of 0 to 1 using the Min-Max scaler along each dimension.
Edge Weights: For edges based on chemical perturbations [4, 6], we assign the edge weight of 1. We also compute cosine similarity for nodes if they are in the same feature space (such as two cell morphology/gene expression profiles, or Morgan fingerprints). To avoid noisy edges from computations, we apply a 0.8 threshold for cosine similarity, and additionally explicitly enforce 99.5% sparsity by selecting top similar edges.
Encoder and Decoder: We use the Graph Isomorphism Model (GIN) [58] as the molecule encoder. All molecules on the context graph are used to pretrain the encoder. Because we extract feature vectors as the decoding targets in different modalities, we could efficiently use Multi-Layer Perception (MLP) as modality decoders. In each training batch, random walks start from the molecule node to extract the walk path. Then, decoders are pretrained to reconstruct corresponding node features from nodes over the path. More details are in appendix B.
@@ 6 Experiments
We focus on three research questions (RQs) regarding InfoAlign’s representation for molecular property prediction, molecule-morphology matching, and performance analysis.
Table 2: Results on ChEMBL2K and Broad6K. We report average AUC (Avg.), as well as the percentage of tasks achieving AUC above 80%, 85%, and 90%. We highlight the best and second best mean. We also highlight the row of the best method in each category.
Dataset ChEMBL2k (AUC $\uparrow$) Broad6k (AUC $\uparrow$) (# Molecule / # Task) (2355 / 41) (6567 / 32) Method Avg. >80% >85% >90% Avg. >80% >85% >90% Morgan Fingerprints MLP 76.8±2.2 48.8±3.9 34.6±6.3 21.9±5.7 63.3±0.3 6.3±0.0 4.4±1.7 3.1±0.0 RF 54.7±0.7 0.0±0.0 0.0±0.0 0.0±0.0 55.5±0.1 0.0±0.0 0.0±0.0 0.0±0.0 GP 51.0±0.0 0.0±0.0 0.0±0.0 0.0±0.0 50.6±0.0 0.0±0.0 0.0±0.0 0.0±0.0 Pretrained GNN AttrMask [18] 73.9±0.5 46.8±2.7 31.2±4.4 14.6±1.7 59.8±0.2 3.1±0.0 3.1±0.0 3.1±0.0 ContextPred [18] 77.0±0.5 55.1±1.3 34.1±4.6 14.6±1.7 60.0±0.2 7.5±1.7 3.1±0.0 3.1±0.0 EdgePred [18] 75.6±0.5 54.2±4.0 34.6±7.2 12.2±2.4 59.9±0.2 3.1±0.0 3.1±0.0 3.1±0.0 GraphCL [60] 75.6±1.6 46.8±7.6 32.2±6.8 18.0±3.7 67.2±0.5 15.6±3.1 3.1±0.0 3.1±0.0 GROVER [44] 73.3±1.4 38.5±2.0 22.4±3.6 14.6±2.4 66.2±0.1 15.6±0.0 3.8±1.4 3.1±0.0 JOAO [60] 75.1±1.0 47.8±5.1 33.7±2.0 19.0±3.2 67.3±0.4 12.5±0.0 3.8±1.4 3.1±0.0 MGSSL [62] 75.1±1.1 39.0±4.6 29.3±3.0 10.3±3.2 66.9±0.5 13.8±2.8 3.1±0.0 3.1±0.0 GraphLoG Xu et al. 73.5±0.7 41.9±2.0 29.3±3.4 15.6±2.8 62.9±0.4 4.4±1.7 0.0±0.0 0.0±0.0 GraphMAE [17] 74.7±0.1 33.2±1.3 27.8±1.3 12.2±1.7 66.8±0.3 14.4±1.7 3.1±0.0 3.1±0.0 DSLA [23] 69.3±1.0 23.9±4.7 14.6±5.5 6.8±1.1 63.3±0.3 6.3±0.0 3.1±0.0 3.1±0.0 UniMol [63] 76.8±0.4 46.8±2.0 33.7±1.1 24.9±2.0 65.4±0.1 7.5±1.7 3.1±0.0 3.1±0.0 Pretrained Chemical Language Models Roberta [31] 74.7±1.9 46.3±3.4 35.1±4.4 22.9±1.3 59.8±0.7 5.0±1.7 3.1±0.0 3.1±0.0 GPT2 [31] 71.0±3.4 31.2±11.2 20.0±9.4 7.3±6.9 60.6±0.3 7.5±1.7 1.9±1.7 1.9±1.7 MolT5 [10] 69.9±0.8 32.2±2.0 21.0±4.1 8.8±1.3 56.4±0.8 3.8±1.4 2.5±1.4 2.5±1.4 ChemGPT [12] 65.0±1.1 16.1±2.8 11.2±3.3 5.4±1.1 55.1±0.9 3.1±0.0 3.1±0.0 1.3±1.7 Cell Morphology MLP 64.3±2.4 15.6±6.6 8.3±3.7 4.9±3.9 51.9±1.0 0.0±0.0 0.0±0.0 0.0±0.0 RF 55.9±0.7 3.9±1.3 3.9±1.3 2.4±0.0 55.3±0.1 0.0±0.0 0.0±0.0 0.0±0.0 GP 50.1±0.0 0.0±0.0 0.0±0.0 0.0±0.0 54.7±0.0 0.0±0.0 0.0±0.0 0.0±0.0 Gene Expression MLP 56.1±1.1 5.1±1.4 3.4±1.3 3.4±1.3 56.9±1.4 1.9±1.7 1.9±1.7 1.9±1.7 RF 52.8±0.3 0.0±0.0 0.0±0.0 0.0±0.0 55.2±0.2 0.0±0.0 0.0±0.0 0.0±0.0 GP Run out of time 50.1±0.0 0.0±0.0 0.0±0.0 0.0±0.0 Multi-modal Alignment CLOOME 66.7±1.8 26.8±4.6 16.1±3.7 10.7±5.1 61.7±0.4 3.1±0.0 3.1±0.0 0.0±0.0 InfoCore (GE) 79.3±0.9 62.4±2.8 46.3±3.0 30.3±2.2 60.2±0.2 3.1±0.0 0.0±0.0 0.0±0.0 InfoCore (CP) 73.8±2.0 37.6±9.2 26.3±4.7 10.7±4.1 61.1±0.2 6.3±0.0 3.1±0.0 0.0±0.0 InfoAlign (Ours) 81.3±0.6 66.3±2.7 49.3±2.7 35.1±3.7 70.0±0.1 18.8±2.2 3.1±0.0 3.1±0.0
6.1 RQ1: Molecular Property Prediction
@@ 6.1.1 Experimental Setting
Dataset and Evaluation: We select datasets for important tasks including activity classification for various assays in ChEMBL2K [13] and Broad6K [33], drug toxicity classification using ToxCast [42], and absorption, distribution, metabolism, and excretion (ADME) regression using Biogen3K [11]. The dataset statistics are in Table 1, covering 685 tasks. We apply scaffold-splitting for all datasets. We follow [19] for the ToxCast dataset, and a 0.6:0.15:0.25 ratio for training, validation, and test sets for other datasets. We use the Area under the curve (AUC) for classification and mean absolute error (MAE) for regression. Mean and standard deviations are reported from ten runs.
@@
Baseline: We include 27 baselines across six categories: (1) three molecular fingerprint (FP)-based methods [43]; (2) eleven pretrained GNNs; (3) four pretrained chemical language models; (4,5) six methods based on cell morphology and gene expression values from cells treated with each molecule; (6) CLOOME [46] and InfoCORE [54] for multi-modal alignment using structure, morphology, and gene expression data. We use MLPs, Random Forests (RF), and Gaussian Processes (GP) for methods in categories (1,4,5). We fine-tune MLPs on various representation learning approaches for predicting molecular properties. Setting details and all results are in Sections C.1 and C.3.
Table 3: Results on ToxCast and Biogen3K. We report the average AUC and the percentage of AUC above 80% on ToxCast, and regression MAE (scaled by × 100) for Biogen3K. We highlight the best and second best mean. We also highlight the row of the best method in each category.
Dataset ToxCast (AUC $\uparrow$) Biogen3K (MAE ×100$\times$ 100$\downarrow$ 100 ↓ ) (# Molecule / # Task) (8576 / 617) (3521 / 6) Method Avg. >80 % Avg. hPPB rPPB RLM HLM ER Solubility Morgan Fingerprints MLP 57.6±1.0 1.6±0.3 66.2±2.4 66.1±2.6 56.8±2.3 56.5±4.2 74.6±6.2 73.7±7.3 69.5±3.0 RF 52.3±0.1 0.2±0.1 52.8±0.2 44.2±0.1 44.2±0.1 42.0±0.2 67.7±0.7 66.9±0.9 51.6±0.1 GP Run out of Time 60.0±0.0 51.3±0.0 59.5±0.0 49.7±0.0 68.8±0.0 69.3±0.0 61.6±0.0 Pretrained GNN AttrMask [18] 63.1±0.8 3.2±1.2 67.3±0.3 82.4±1.1 49.8±0.7 51.7±1.0 57.9±0.6 62.6±0.5 99.1±1.2 ContextPred [18] 63.0±0.6 3.3±1.3 68.5±0.9 85.0±7.9 49.7±0.4 55.1±2.7 61.4±1.8 63.1±0.5 96.5±3.7 EdgePred [18] 63.5±1.1 4.8±3.0 67.8±0.9 81.2±10.2 48.0±0.5 53.5±2.8 62.2±1.8 62.9±0.7 99.1±6.9 GraphCL [60] 52.2±0.2 0.5±0.3 53.9±0.6 43.8±0.3 45.4±0.6 40.6±0.5 76.7±1.0 67.1±2.2 49.6±0.3 GROVER [44] 53.1±0.4 0.5±0.1 54.9±1.6 44.5±0.4 46.5±0.7 41.7±0.6 73.2±5.7 71.0±4.3 52.6±0.3 JOAO [61] 52.3±0.2 0.4±0.1 55.0±0.8 44.5±0.5 47.6±0.5 40.6±0.2 74.3±2.8 71.5±2.6 51.4±0.6 MGSSL [62] 64.2±0.2 4.0±0.4 53.2±0.3 44.8±0.6 49.7±0.3 41.5±0.2 65.6±1.8 64.6±0.5 52.7±0.5 GraphLoG [59] 58.6±0.4 2.5±0.3 56.9±0.4 49.3±0.3 54.8±0.5 42.6±0.3 66.8±1.7 69.0±1.3 58.8±0.5 GraphMAE [17] 53.3±0.1 0.6±0.1 52.8±0.8 43.3±0.9 51.2±0.8 40.9±0.3 64.4±2.7 65.9±3.8 50.9±1.4 DSLA [23] 57.8±0.5 0.7±0.1 57.9±0.7 50.4±0.7 53.6±1.7 43.3±0.9 68.6±1.2 70.8±2.0 60.9±0.6 UniMol [63] 64.6±0.2 4.8±1.0 55.8±2.8 50.1±5.2 49.9±5.6 43.6±1.1 65.4±4.9 65.8±1.2 59.9±6.6 Pretrained Chemical Language Models Roberta [31] 64.2±0.8 3.1±1.8 69.0±2.6 71.4±14.5 65.1±19.2 63.7±24.6 67.5±5.2 69.9±4.9 76.7±13.2 GPT2 [31] 61.5±1.1 2.4±0.6 74.0±8.5 65.4±12.9 73.1±20.8 54.1±12.9 83.2±21.5 86.1±19.8 81.8±25.5 MolT5 [10] 64.7±0.9 3.6±1.1 65.1±0.5 76.7±2.1 55.9±1.1 49.2±1.0 70.3±0.8 73.1±1.0 65.3±1.7 ChemGPT [12] Token Error 75.7±8.5 59.5±7.3 88.8±32.3 76.1±11.8 84.0±20.6 77.2±8.5 68.6±7.1 Multi-modal Alignment CLOOME 54.2±0.9 0.9±0.2 64.3±0.4 65.2±1.5 56.9±0.8 44.2±0.8 70.7±0.4 73.6±0.8 75.0±2.1 InfoCORE (GE) 65.3±0.2 5.4±1.7 69.9±1.2 79.9±3.6 51.6±1.8 51.3±2.1 78.6±0.3 77.8±1.9 80.3±0.9 InfoCORE (CP) 62.4±0.4 1.3±0.5 71.0±0.6 74.5±4.9 53.5±0.7 53.6±2.1 80.8±1.5 79.4±3.4 84.4±1.0 InfoAlign (Ours) 66.4±1.1 6.6±1.6 49.4±0.2 39.7±0.4 39.2±0.3 40.5±0.6 66.7±1.7 62.0±1.5 48.4±0.6
6.1.2 Results and Analysis
We present results across various assays in Tables 2 and 3 and Figure 3. Key observations include:
(1) Molecular structures are superior compared to cell morphology and gene expression features for predicting various molecular assays. This is likely because the datasets and tasks we selected fundamentally involve predicting the binding affinity of a molecule to a protein [13]; furthermore, in these datasets, molecules with activity in a given assay tend to have highly related structures, rather than representing two or more structurally distinct classes of molecules with activity; together this implies that molecular structure alone will tend to yield strong results. When comparing the three popular structure-based representation approaches, no single method outperforms the others across all four datasets. Pretrained GNNs generally perform better than fingerprint-based methods and pretrained chemical language models, thanks to recent advancements. However, continued efforts in universal structural representation are still necessary.
(2) Cell morphology and gene expression features may complement molecular structures, yielding more generalizable representations. As shown in Figure 3, cell morphology and gene expression outperform molecular structure in approximately 20% and 10% of tasks on the ChEMBL2K and Broad6K datasets, respectively. This suggests that incorporating cell context into representation learning would be beneficial. That said, existing multi-modal baselines (InfoCORE, CLOOME) only outperform molecular structure-based approaches on ChEMBL2K and ToxCast, as they do not construct molecular representations holistically by using all cell-related modalities.
(3) InfoAlign achieves the best average performance on all tasks compared to 27 baselines. The improvements from InfoAlign range from 2.5% to 6.4% on average across four datasets compared to the second-best method. These gains are more significant when using the 80% AUC threshold on classification datasets. While InfoCORE (GE) performs best among baselines on the ChEMBL2K and ToxCast datasets, it struggles to align molecular representations with more than two modalities and sometimes leads to negative transfer, as seen in Broad6K and Biogen3K.